# Flow Matching for Collaborative Filtering

### v1

@@ Abstract.
Generative models have shown great promise in collaborative filtering by capturing the underlying distribution of user interests and preferences. However, existing approaches struggle with inaccurate posterior approximations and misalignment with the discrete nature of recommendation data, limiting their expressiveness and real-world performance. To address these limitations, we propose FlowCF, a novel flow-based recommendation system leveraging flow matching for collaborative filtering. We tailor flow matching to the unique challenges in recommendation through two key innovations: (1) a behavior-guided prior that aligns with user behavior patterns to handle the sparse and heterogeneous user-item interactions, and (2) a discrete flow framework to preserve the binary nature of implicit feedback while maintaining the benefits of flow matching, such as stable training and efficient inference. Extensive experiments demonstrate that FlowCF achieves state-of-the-art recommendation accuracy across various datasets with the fastest inference speed, making it a compelling approach for real-world recommender systems.
@@
ccs: Information systems Recommender systems
@@ 1. Introduction
Collaborative filtering (CF) – as the foundational component of model-based recommender systems (schafer2007collaborative; koren2021advances) – aims to recommend items to users by leveraging the preferences of similar users based on their past interactions with items. One exciting thread is the rise of generative collaborative filtering, ranging from variational autoencoders (VAEs) (kingma2013auto) to more advanced diffusion models (ho2020denoising; song2020score). These generative CF approaches model the distribution of user preferences to generate diverse and realistic interaction patterns, thus better handling challenges like noisy data (wang2021denoising) and cold-start scenarios (chen2022generative), in comparison with non-generative methods that focus on learning user and item representations for prediction.
@@ 2.1. Generative Collaborative Filtering
We denote the set of users as $\mathcal{U}=\{u_{i}\}_{i=1}^{|\mathcal{U}|}$, the set of items as =$\mathcal{I}=\{i_{i}\}_{i=1}^{|\mathcal{I}|}$, and the user-item interaction matrix as $\bar{X}\in\{0,1\}^{|\mathcal{U}|\times|\mathcal{I}|}$. Recommendation systems aim to predict the missing entries in $\bar{X}$ effectively, uncovering unobserved user preferences and potential interactions. CF approaches this problem by recommending items to users based on the interaction patterns of other users with similar interests and preferences.
@@ 2.2. Flow Matching
(2) Find $\text{Find }v_{t}(x)\text{ generating the probability path }p_{t},$
@@ 3.2. Behavior-Guided Prior (RQ1)
(9) $\mathrm{f}_{i}=\frac{\text{number of interactions for item }i}{\text{total number of users }|\mathcal{U}|}.$
@@ 3.3. Discrete Flow Framework (RQ2)
Thus, despite the discretization, the discrete flow framework maintains a theoretical connection to continuous state space.
@@ 3.4. Training FlowCF (RQ3)
1:Trained flow model $f_{\theta}$ with parameters $\theta$, observed user interactions $X$, number of discretization steps $N$
@@
3.5. Inference with FlowCF (RQ4)
During the inference process, the observed user interactions $X$ can be treated as a noisy interaction matrix, analogous to an interpolation $X_{t}$ between prior sample $X_{0}$ from the behavior-guided prior and the ground truth interactions $X_{1}$. To initiate the inference process, a starting sampling step needs to be selected. Since the starting point $X$ is not purely noise but contains meaningful information about user historical behavior, it is unnecessary to begin at $t=0$. Instead, the inference process can start from a later step, significantly reducing the number of sampling steps required. To preserve the binary nature of the interactions $X_{t}$ throughout the inference process, FlowCF employs the following update rule to compute $X_{t+\frac{1}{N}}$ at the next step of $t$:
@@ 4.1. Experimental Setup
Datasets. To evaluate our proposed model, we conduct our experiments on three public datasets in real-world scenarios: MovieLens-1M (harper2015movielens), MovieLens-20M, and Amazon-Beauty (mcauley2015image). These datasets vary in scale and density, which allow us to evaluate the performance of the models across different levels of sparsity. Although they consist of explicit ratings, we have intentionally chosen them to investigate the performance of learning from implicit feedback. To convert explicit ratings into implicit feedback, we binarize the ratings by treating ratings of four and above as positive interactions and discarding ratings ¡ 4. We split the dataset into training, validation, and test sets at a ratio of 8:1:1. To be consistent with the preprocessing in previous studies (liang2018variational), we exclude users and items with fewer than five interactions from the datasets. Table 2 provides a summary of the statistics of the datasets after preprocessing.
@@
Baselines. We compare FlowCF with several baselines, including non-generative methods like MF-BPR (rendle2012bpr), LightGCN (he2020lightgcn), and SGL (wujc2021self); DAE-based methods like CDAE (wu2016collaborative) and Mult-DAE (liang2018variational); VAE-based methods like Mult-VAE (liang2018variational), MacridVAE (ma2019learning), and RecVAE (shenbin2020recvae); and the diffusion-based method DiffRec (wang2023diffusion). The detailed descriptions of baselines are in Appendix B.
@@
Implementation Details. We use 9 discretization steps for ML-1M and Amazon-Beauty, and 50 for ML-20M. The number of sampling steps is fixed at 2. Further details are provided in Appendix C. The code is available at https://github.com/chengkai-liu/FlowCF.
@@ 4.3. Ablation and Sensitivity Study
Effect of Priors and Flows. To study the effect of different prior distributions on both continuous flow-based recommender (denoted as CFlowCF, further details in Appendix A) and discrete flow-based FlowCF, we compare uniform distribution between 0 and 1, Gaussian distribution $\mathcal{N}(0,I)$, and our behavior-guided prior distribution. For discrete flow-based FlowCF, we discretize these priors to probability mass functions using a Bernoulli distribution and also include a random binary initialization (0 or 1) as a baseline.
@@
These findings indicate that although the discrete flow framework improves recommendation performance, it is more sensitive to the choice of prior. This sensitivity arises because it operates in a non-differentiable discrete state space compared to smooth differentiable continuous space, necessitating the customized behavior-guided prior for effective training. Additionally, we observe that the convergence speed is significantly slower when using priors other than the behavior-guided prior for FlowCF. Overall, these results highlight the importance of incorporating user behavior patterns into the prior distribution to effectively model implicit feedback and enhance recommendation performance in flow matching recommenders.
@@ 5.2. Generative Models
Diffusion Models  (sohl2015deep; song2019generative; ho2020denoising; song2020score) generate data by iteratively refining samples from Gaussian noise to meaningful outputs. These models have achieved strong results across domains, including images (ho2020denoising; rombach2022high), molecules (xu2022geodiff; jing2022torsional), and recommendation systems (lin2024survey; li2023diffurec; yang2024generate).
Flow Matching  (lipman2022flow; liu2022flow; albergo2022building) is a flow-based method using continuous normalizing flows (chen2018neural) as alternatives to diffusion models. Conditional Flow Matching (lipman2022flow; tong2023improving) directly learns a vector field that defines the probability flow from a source distribution to the target distribution, conditioned on individual data points. Flow matching offers faster inference and improved efficiency and has been applied to various generative tasks such as images (esser2403scaling), videos (polyak2024movie), and proteins (bose2023se). However, its potential in recommender systems remains unexplored, presenting opportunities for future research.
@@ 6. Conclusion
Appendix A Continuous Flow-Based CFlowCF
@@
Appendix B Details of the Baselines
@@
Appendix C Implementation Details

### v2

@@ Abstract.
Generative models have shown great promise in collaborative filtering by capturing the underlying distribution of user interests and preferences. However, existing approaches struggle with inaccurate posterior approximations and misalignment with the discrete nature of recommendation data, limiting their expressiveness and real-world performance. To address these limitations, we propose FlowCF, a novel flow-based recommendation system leveraging flow matching for collaborative filtering. We tailor flow matching to the unique challenges in recommendation through two key innovations: (1) a behavior-guided prior that aligns with user behavior patterns to handle the sparse and heterogeneous user-item interactions, and (2) a discrete flow framework to preserve the binary nature of implicit feedback while maintaining the benefits of flow matching, such as stable training and efficient inference. Extensive experiments demonstrate that FlowCF achieves state-of-the-art recommendation accuracy across various datasets with the fastest inference speed, making it a compelling approach for real-world recommender systems. The code is available at https://github.com/chengkai-liu/FlowCF.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736967isbn: 979-8-4007-1454-2/2025/08ccs: Information systems Recommender systems
@@ 1. Introduction
Collaborative filtering (CF) (schafer2007collaborative; koren2021advances) – as the foundational component of model-based recommender systems (liu2024mamba4rec; liu2024behavior) – aims to recommend items to users by leveraging the preferences of similar users based on their past interactions with items. One exciting thread is the rise of generative collaborative filtering, ranging from variational autoencoders (VAEs) (kingma2013auto) to more advanced diffusion models (ho2020denoising; song2020score). These generative CF approaches model the distribution of user preferences to generate diverse and realistic interaction patterns, thus better handling challenges like noisy data (wang2021denoising) and cold-start scenarios (chen2022generative), in comparison with non-generative methods that focus on learning user and item representations for prediction (he2017neural; he2020lightgcn).
@@ 2.1. Generative Collaborative Filtering
We denote the set of users as $\mathcal{U}=\{u_{i}\}_{i=1}^{|\mathcal{U}|}$, the set of items as $\mathcal{I}=\{i_{i}\}_{i=1}^{|\mathcal{I}|}$, and the user-item interaction matrix as $\bar{X}\in\{0,1\}^{|\mathcal{U}|\times|\mathcal{I}|}$. Recommendation systems aim to predict the missing entries in $\bar{X}$ effectively, uncovering unobserved user preferences and potential interactions. CF approaches this problem by recommending items to users based on the interaction patterns of other users with similar interests and preferences.
@@ 2.2. Flow Matching
(2) $\text{Find }v_{t}(x)\text{ generating the probability path }p_{t},$
@@ 3.2. Behavior-Guided Prior (RQ1)
[…]
@@ 3.3. Discrete Flow Framework (RQ2)
Therefore, despite the use of discrete variables, the discrete flow framework maintains a theoretical alignment with continuous interpolation in expectation.
@@ 3.4. Training FlowCF (RQ3)
1:Trained flow model $f_{\theta}$ with parameters $\theta$, observed user interactions $X$, number of discretization steps NN
@@
3.5. Inference with FlowCF (RQ4)
During the inference process, the observed user interactions $X$ can be treated as a noisy interaction matrix, analogous to an interpolation $X_{t}$ between prior sample $X_{0}$ from the behavior-guided prior and the ground truth interactions $X_{1}$. To initiate the inference process, a starting sampling step needs to be selected. Since the starting point $X$ is not purely noise but contains meaningful information about user historical behavior, it is unnecessary to begin at $t=0$. Instead, the inference process can start from a later step, significantly reducing the number of sampling steps required. To preserve the binary nature of the interactions $X_{t}$ throughout the inference process, FlowCF employs the following update rule, where each update corresponds to an element-wise argmax over the predicted field, to compute $X_{t+\frac{1}{N}}$ at the next step of $t$:
@@ 4.1. Experimental Setup
Datasets. To evaluate our proposed model, we conduct our experiments on three public datasets in real-world scenarios: MovieLens-1M (harper2015movielens), MovieLens-20M, and Amazon-Beauty (mcauley2015image). These datasets vary in scale and density, which allows us to evaluate the performance of the models across different levels of sparsity. Although they consist of explicit ratings, we have intentionally chosen them to investigate the performance of learning from implicit feedback. To convert explicit ratings into implicit feedback, we binarize the ratings by treating ratings of four and above as positive interactions and discarding ratings ¡ 4. We split the dataset into training, validation, and test sets at a ratio of 8:1:1. To be consistent with the preprocessing in previous studies (liang2018variational), we exclude users and items with fewer than five interactions from the datasets. Table 2 provides a summary of the statistics of the datasets after preprocessing.
@@
Baselines. We compare FlowCF with several baselines, including non-generative methods like MF-BPR (rendle2012bpr), LightGCN (he2020lightgcn), and SGL (wujc2021self); DAE-based methods like CDAE (wu2016collaborative) and Mult-DAE (liang2018variational); VAE-based methods like Mult-VAE (liang2018variational), MacridVAE (ma2019learning), and RecVAE (shenbin2020recvae); and the diffusion-based method DiffRec (wang2023diffusion). The detailed descriptions of baselines are in Appendix C.
@@
Implementation Details. We use 9 discretization steps for ML-1M and Amazon-Beauty, and 50 for ML-20M. The number of sampling steps is fixed at 2. Further details are provided in Appendix D.
@@ 4.3. Ablation and Sensitivity Study
Effect of Priors and Flows. To study the effect of different prior distributions on both continuous flow-based recommender (denoted as CFlowCF, further details in Appendix B) and discrete flow-based FlowCF, we compare uniform distribution between 0 and 1, Gaussian distribution $\mathcal{N}(0,I)$, and our behavior-guided prior distribution. For discrete flow-based FlowCF, we discretize these priors to probability mass functions using a Bernoulli distribution and also include a random binary initialization (0 or 1) as a baseline.
@@
These results indicate that although the discrete flow framework improves recommendation performance, it is more sensitive to the choice of prior. This sensitivity arises from the inherent discrete nature and the density of interaction data. In discrete flow framework, the sample $X_{0}$ from prior is binary (only 0s and 1s). To learn effective transformations, the density of positives (1s) in $X_{0}$ should closely match the true density of the dataset. Dense samples from uniform or Gaussian priors introduce many false positives, hindering the model to recover correct patterns. In contrast, CFlowCF uses non-binary continuous-valued samples, which are not affected by the density of positives in discrete modeling. This explains its lower sensitivity to prior design compared to FlowCF. Consequently, a customized behavior-guided prior for effective training. Additionally, we observe that the convergence speed is significantly slower when using priors other than the behavior-guided prior for FlowCF. Overall, these results highlight the importance of incorporating user behavior patterns into the prior distribution to effectively model implicit feedback and enhance recommendation performance in flow matching recommenders.
@@ 5.2. Generative Models
Diffusion Models  (sohl2015deep; song2019generative; ho2020denoising; song2020score) generate data by iteratively refining samples from Gaussian noise to meaningful outputs. These models have achieved strong results across domains, including images (ho2020denoising; rombach2022high), molecules (xu2022geodiff; jing2022torsional), material science (xie2021crystal; luo2023towards) and recommendation systems (lin2024survey; li2023diffurec; yang2024generate).
Flow Matching  (lipman2022flow; liu2022flow; albergo2022building) is a flow-based method using continuous normalizing flows (chen2018neural) as alternatives to diffusion models. Conditional Flow Matching (lipman2022flow; tong2023improving) directly learns a vector field that defines the probability flow from a source distribution to the target distribution, conditioned on individual data points. Additionally, recent advances have also extended the flow matching framework to discrete state spaces (campbell2024generative; gat2024discrete). Flow matching offers faster inference and improved efficiency and has been applied to various generative tasks such as images (esser2403scaling), videos (polyak2024movie), and proteins (bose2023se). However, its potential in recommender systems remains unexplored, presenting opportunities for future research.
@@ 6. Conclusion
Acknowledgements.
We thank the anonymous reviewers and area chairs for their valuable feedback and suggestions. This research was supported in part by Yale AI Engineering Research Seed Grants, Yale Office of the Provost, and Amazon Research Award to the Yale University team.
Appendix A Appendix
Algorithm 3 Inference with CFlowCF
1:Trained flow model $f_{\theta}$ with parameters $\theta$, observed user interactions $X$, number of discretization steps NN
2:Set the starting step $s$
3:Initialize $X_{t}\leftarrow X$
4:for $i=s\text{ to }N-2$ do
5:  Set the current step $t\leftarrow t_{i}$
6:  Predict $\hat{X}=f_{\theta}(X_{t},t)$
7:  Compute $v_{t}\leftarrow(\hat{X}-X_{t})/(1-t)$
8:  Update $X_{t}\leftarrow X_{t}+v_{t}/N$
9:end for
10:Predict $\hat{X}=f_{\theta}(X_{t},t)$ in the last step $t=t_{N-1}$
11:Predicted interactions $\hat{X}$
Appendix B Continuous Flow-Based CFlowCF
@@
Appendix C Details of the Baselines
@@
Appendix D Implementation Details